APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday August 16, 2010 (MJD 55425) ---=== OBSERVING TEAM ===--- Steph Snedden Kaike Pan Paul Harding ---=== OBSERVING PLAN ===--- Guider tests, possible MARVELS science ---=== OBSERVING SUMMARY ===--- Most of the night was lost to humidity. When we were closing, Paul ran some guider focus test, and moved rear lens about 3/8" from the infinite. We tested point model with a BOSS plate. Stars were found on gotoFeild. We observed one MARVELS plate: plate 3703. Stars were found on togoField. We had an alt slip on alt motor #1 at 6 degree. It seemed that 30 lb movement of counterweight makes a visible difference on telescope balance. If so, do we need move counterweight during night to count the difference between full and empty 10 lit dewars? SP1 180L dewar needs to be replaced. ---=== OBSERVING LOG ===--- Afternoon: ---------- All systems and actors were checked. No probelm to report. Paul is running guiderActor from his directory. The version is not checked in yet. Night: ------ 02:35Z high humidity prevents us from opening. 02:45Z started guider focus tests. During the test, it was found that the guider could not be focused with both lenses in infinite. The rear lens had to be 5 knurls away from the infinite (~3/8" on the circumfrence). Guider scale was measured, it is with 1%. The projected scale is about 12.89 um/pix instead of 13 um/pix (by designed). 04:00Z Humidity is about 80%, good enough to open enclosure. Switch guiderActor to verion:"v1_0_20h" load cartridge 11, plate 4097. 04:20Z slew to the field to test pointing model. stars are in acquiring fibers on goto field. Best focus on this cartridge is 283. STUI Guider GUI on Paul's machine does not work. 05:00Z we had to close for high humidity. 09:20Z humidity and sky both were improving, started ventilation. 09:46Z MARVELS bias and calibs T09:46:15.060.fits BIAS 7.0000000 T09:52:55.980.fits TIO 46877.891 T09:55:01.290.fits THAR 313.75000 T09:59:36.790.fits TIO 47332.891 10:00Z enclosure off. We had a alt slip at 6 degree. It is on alt #1. 10:15Z cart 3, plate 3703 A. SOGTF. cloudy with okay seeing (1.3"). T10:24:08.460.fits STAR 34.083668 Possible broken fiber: 93/94 11:19Z MARVELS calibs T11:17:37.970.fits TIO 48125.109 T11:19:43.280.fits THAR 317.14374 T11:24:17.590.fits TIO 48158.555 11:20Z Mount cart 11, a BOSS cartridge. Moved down counterweight by 30 lb. It seemed that 30 lb did make difference in telescope balance. We did not have alt slip when we went to 6 degree. 11:30Z enclosure on. 11:45Z endNight finished. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 10:44Z 3 0 169.75 0.0000 49.42 0.0000 -28.30 0.0000 1.000420 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 10:44Z 3 0 1.00042 -4148 -1973 277 170 49.4 14.0 10 198 -- 1.3 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 01:43Z 60 51 5 192 (SSW) -9999 - 10:45Z 57 44 10 198 (SSW) -9999 - ---=== TELESCOPE STATUS ===--- 01:00Z - Began ventilation. Doors and louvers opened, fans on. 04:00Z - Enclosure off. 05:00Z - Enclosure on, doors and louvers closed, fans off. 09:20Z - Began ventilation. Doors and louvers opened, fans on. 10:00Z - Enclosure off. 11:30Z - Enclosure on, doors and louvers closed, fans off No interlocks bypassed. Status at: 11:50Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 11 Counterweights at: 240 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: SP1 Spectro 264 lbs, 18.8 psi SP2 33 lb, 20.6 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 MCP: v6_6_5 TPM: tpm_v3_5_0 STUI: 1.0b2 2010-07-21 alerts: trunk+svn114742 apo: trunk+svn114701 boss: trunk+svn115542M boss.daq: 1.2.14 boss.specMech: Offline gcamera: trunk+svn115486M guider: ** FAILED ** (running paul's version, not checked in). hub: Tag: UNTAGGED_REVISION 502:504M mcp: v6_6_5 mcp.plc: 25 platedb: trunk+svn111739 sop: v1_1_3 sos: ** FAILED ** tcc: 2.13.0 Pidlspec2d (sos3) --- v5_4_12 plate-mapper3 ------ NOCVS:v5_0_7rocedure: v3_06 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.5090 -5.5100 Axial B 16.0950 16.0940 Axial C 15.9220 15.9220 Trans D 11.0070 11.0100 Lateral E 10.6553 10.6553 Lateral F 12.2301 12.2301 GALILS Commanded: 10600. -250. 5800. 10250. 13050. 12650. Actual: 10600. -250. 5800. 10250. 13050. 12650. SETMIR VALUES PriDesOrient: 0.00 -13.67 10.53 574.80 742.37 PriOrient: 0.00 -13.67 10.53 574.80 742.37 SECONDARY: ---------- Focus: 0.00 Air Temp.: 15.6 Alt.: 29.999997 MIGS TONIGHT NOMINAL Axial A -7.8330 -7.8320 Axial B -0.0440 -0.0410 Axial C 13.8180 13.8180 Trans D 14.1830 14.1840 GALILS Commanded: 1585369. 1566742. 1598868. -1450. -50. Actual: 1585350. 1566700. 1598900. -1450. -50. SETMIR VALUES SecDesOrient: 1257.00 -9.00 -4.00 -58.00 23.67 SecOrient: 1257.01 -8.98 -3.99 -58.08 23.30 ---=== PROBLEMS IN DETAIL ===---